Write a quadratic equation in the variable x having the given numbers as solutions.Type the equation in standard form.
Solution 9 only solution.

Answer by ewatrrr(24785)

y = (x-9)^2
(x-9)(x-9) = 0
x^2 - 18x + 81 = 0
F First terms
O Outside terms
I Inside terms
L Last terms